OpenStudy (anonymous):

Solve the equation. |k + 7| = 3 A. {–4, 4} B. {–10, –4} C. {all real numbers greater than or equal to –10 and less than or equal to – 4} D. {–4, 10}

OpenStudy (anonymous):

@mikayla16

OpenStudy (anonymous):

k+7=3 k=3-7 k=-4 k+7=-3 k=-3-7 k=-10 So B

OpenStudy (anonymous):

d

OpenStudy (anonymous):

Equation to Solve |k+7|=3 Solution Steps 1. Remove Absolute Value Forked Equation 1 k+7=3 Forked Equation 2 −(k+7)=3 Now Solving Equation 1 k+7=3 2. Move Constants to Right k=3−7 3. Add Constants k=-4 End of Equation 1 Now Solving Equation 2 −(k+7)=3 4. Negate -k−7=3 5. Move Constants to Right -k=3+7 6. Add Constants -k=10 7. Negate Both Sides k=−10 End of Equation 2 Plug candidate answers into original equation... Answer 1: -4 Answer 2: -10